OpenStudy (anonymous):

if csc^-1x = sin^-1(1/x) then is sin^-1 = csc^-1(1/x) ???

OpenStudy (anonymous):

please help me!!

OpenStudy (anonymous):

is that a question ?

OpenStudy (anonymous):

yes i was wondering if this is right

OpenStudy (anonymous):

i want to rewrite y=csc^-1x to sin but i dont know..

OpenStudy (anonymous):

kk ! I'll ask you a question : what is the domain of the function \[\csc^{-1}\]

OpenStudy (anonymous):

x≤-1 or x≥1

OpenStudy (anonymous):

now if \[x \ge 1=====>\frac{1}{x}\text{ ?????? } \] ?!

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \ \div \ x≥1\end{matrix}\right)\]

OpenStudy (anonymous):

Now I'm lost ! what did you prove ! this \[\csc^{-1}(x)=\sin^{-1}(\frac{1}{x})\] ?!

OpenStudy (anonymous):

i just want to know if this is correct but i dont know how to solve when sin^-1(x)

OpenStudy (anonymous):

Sorry im confusing you:(

OpenStudy (anonymous):

No !No ! problem at all ! let's discuss about that ! what you diid is correct...If \[x \ge 1 \text{ or } x \le -1 \]

OpenStudy (anonymous):

it can be any numbers right? x≥1 or x≤−1

OpenStudy (anonymous):

Yes ! So, keep this in your "japanese " mind if we have to write i(or type) \[\csc^{-1}(***)\text{ then (***) should be } *** \ge 1 or ***\]

OpenStudy (anonymous):

*** less than -1 ! now let's return back to your question !

OpenStudy (anonymous):

hmmm...? im lost i tried this out is it right?

OpenStudy (anonymous):

It seems correct ! But It's not ! You did the biggest mistake of your life ! that what I was saying ! You should be careful ! if \[\csc^{-1}(x)=\sin^{-1}(\frac{1}{x})\] then \[x \le-1 \text{ or }x \ge 1 \] we're asked :if \[\csc^{-1}(\frac{1}{x})=\sin^{-1}(x)\]

OpenStudy (anonymous):

then \[\frac{1}{x}\text{ should also satisfy } \frac{1}{x} \le -1 \text{ or } \frac{1}{x} \ge 1 \]

OpenStudy (anonymous):

is that possible ?

OpenStudy (anonymous):

@lsugano

OpenStudy (anonymous):

.......i am so lost :( okay ill try to make it clear again and write it on my tablet and if possible please checkk!!

OpenStudy (anonymous):

kk ! you don't know yet if the statement : if csc^-1x = sin^-1(1/x) then sin^-1(x) = csc^-1(1/x) is correct or not !

OpenStudy (anonymous):

yes! so i was wondering if this is correvct?

OpenStudy (anonymous):

We are looking for an answer together :) :) !

OpenStudy (anonymous):

so is this statement true?

OpenStudy (anonymous):

both !

OpenStudy (anonymous):

True if x=1 false if x=/= 1

OpenStudy (anonymous):

false if x is not 1 you mean?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

so what if it was more than or less than 1?

OpenStudy (anonymous):

for all numbers ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it will be false certainly !

OpenStudy (anonymous):

can you prove that to me ?!

OpenStudy (anonymous):

idont know how....>< this is confusing me

OpenStudy (anonymous):

:'( :'( ! I've just explained it to you ! all of this is about the domain !

OpenStudy (anonymous):

wait i think i got it! so my main goal is to understand the inverse secant cosecant and cotangent of the functions!

OpenStudy (anonymous):

:) :) use your Tablet ! and prove that to me !

OpenStudy (anonymous):

are you doing that ?

OpenStudy (anonymous):

im almost done with csc!

OpenStudy (anonymous):

kk ! I'm waiting...Take your time !

OpenStudy (anonymous):

i got tan!

OpenStudy (anonymous):

Another Question ?

OpenStudy (anonymous):

imeant cotagent!!

OpenStudy (anonymous):

It wasn't csc^-1 ?!

OpenStudy (anonymous):

im trying to find all three i also found csc^-1

OpenStudy (anonymous):

take a look at this !

OpenStudy (anonymous):

what does the arrows stand for?

OpenStudy (anonymous):

\[\text{statement A} <=====>\text{statement B}\] means statement A is equivalent to Satement B